OpenStudy (anonymous):

cos[inverse tangent 5/12-inverse sin (-3/5)]

zepdrix (zepdrix):

Hey Chels! \(\Large \color{royalblue}{\text{Welcome to OpenStudy! :)}}\)

zepdrix (zepdrix):

Hmmm let's see what we've got here...

zepdrix (zepdrix):

\[\Large \cos\left[\arctan\left(\frac{5}{12}\right)-\arcsin\left(-\frac{3}{5}\right)\right]\] \[\Large \arctan\left(\frac{5}{12}\right)=\phi\qquad\qquad \arcsin\left(-\frac{3}{5}\right)=\theta\] Hmm this one is a bit tricky, we'll have to draw 2 separate reference triangles.

zepdrix (zepdrix):

\[\Large \arcsin\left(-\frac{3}{5}\right)=\theta \qquad\to\qquad \sin \theta=-\frac{3}{5}\]

zepdrix (zepdrix):

Understand how to label the triangle? :o